OpenStudy (anonymous):

If x=14 what is 2x+20 how do you work this equation?

OpenStudy (anonymous):

Put 14 in for x. 2(14)+20

OpenStudy (anonymous):

you plug 14 into the equation

OpenStudy (anonymous):

2(14) = 28+20= 48, 48 is your answer

OpenStudy (anonymous):

2(14)+20=48, 48 is the correct answer.

OpenStudy (anonymous):